OpenStudy (anonymous):

a swimsuit is on sale at 20% discount. if it normally sells for $64.99 what is the sale price ?

OpenStudy (bahrom7893):

64.99-100% x-20% x=64.99*20/100

OpenStudy (bahrom7893):

x=12.99

OpenStudy (bahrom7893):

64.99-12.99=$52

OpenStudy (anonymous):

whars the answer

OpenStudy (bahrom7893):

No 12 bucks is how much they took off, final price is $52

OpenStudy (muhammad_nauman_umair):

yes bahrom7893 u r right discount =64.99*20%=64.99*0.2=12.998 sales price =64.99-12.998=51.992